Water at 60°F is pumped from a reservoir to the top of a mountain through in 6-in. Schedule 120 pipe at an average velocity of 12ft/s. The pipe discharges into the atmosphere at a level 3000ft above the level in the reservoir. The pipeline itself is 4500ft long. Of overall efficiency to the pump and the motor driving it is 70% and the cost of electric energy to the motor is 4 cents per kilowatt hour, what is the hourly energy cost for pumping this water? (In fluid mechanics)
1
Expert's answer
2020-04-02T08:59:48-0400

Mass of water flown per second"=" Volume Flow rate"\times" Density"=Av\times \rho"

Diameter of pipe"=6\ in=0.5\ ft=0.5\times 0.305\ m=0.1525\ m"

Area of pipe"=\frac{\pi d^2}{4}=0.018256 m^2"

Density of water"=1000Kg\/m^3"

Velocity of water"=12ft/sec=0.305\times 12m/sec=3.66m/sec"

.Mass flow rate"(m_r)=0.018256\times3.66\times1000=66.81Kg/sec"

Total energy per second required to reach water upto height of "3000\ ft \ or\ 3000\times 0.305\ m(915\ m)=m_rgh=" "66.81\times 9.81\times 915" "=599.15\ KW"

As per question,motor has efficiency of "70" % and motor power be "P."

So,"\frac{70}{100}\times P=599.15\ KW"

"\implies P=855.932\ KW"

Cost is "4" cents per "KW\ h" .

Hourly cost will be "4\times 855.932=3423.73" cents